the equation of a transverse wave travelling along a string is given by Y=1.8sin(23.8X+ 317t) where X is in meter and Y is in millimeter and is in seconds .The string under tension of 16.3N. Find linear mass density?


1
Expert's answer
2021-02-09T13:33:52-0500

We can find the linear mass density from the formula:


"v=\sqrt{\dfrac{T}{\mu}},""\mu=\dfrac{T}{v^2}."

We can find the speed of the wave from the wave speed equation:


"v=f\lambda."

We can find the frequency from the formula:


"f=\dfrac{\omega}{2\pi}=\dfrac{317\ \dfrac{rad}{s}}{2\pi}=50.45\ Hz."

We can find the wavelength of the wave from the formula:


"\lambda=\dfrac{2\pi}{k}=\dfrac{2\pi}{23.8\ m^{-1}}=0.26\ m."

Finally, we can find the linear mass density:

"\mu=\dfrac{T}{(f\lambda)^2}=\dfrac{16.3\ N}{(50.45\ Hz\cdot0.26\ m)^2}=9.47\cdot10^{-2}\ \dfrac{kg}{m}."